OpenStudy (anonymous):

Which of the following most accurately describes the translation of the graph y = (x + 3)2 to the graph of y = x2 + 3 ? The graph stays the same. left 3, no change up or down up 3, no change left or right up 3 and 3 to the right

OpenStudy (triciaal):

did you view the graphs?

OpenStudy (triciaal):

inside the parenthesis moves the graph left or right added to the graph moves it up or down

OpenStudy (triciaal):

look at the graph how do you get from red to blue?

OpenStudy (anonymous):

up 3 and 2 the right 3 ? @triciaal

OpenStudy (anonymous):

Which of the following MOST accurately describes the translation of the graph y = (x + 1)2 + 4 to the graph of y = (x + 5)2 – 2? down 6 and 5 to the left down 2 and 5 to the left up 4 and 1 to the left

OpenStudy (triciaal):

yes D for the first one

OpenStudy (triciaal):

change in x 5 -1 =4 units moved from x =-1 to x = -5

OpenStudy (triciaal):

did you forget one of the options?

OpenStudy (anonymous):

down 6 and 5 to the left down 2 and 5 to the left up 4 and 1 to the left down 6 and 4 to the left

OpenStudy (anonymous):

so the second one would be down 6 to the left 4 ?

OpenStudy (anonymous):

What happens to the vertex of the graph of y = (x – 4)2 if the equation is changed to y = (x + 4)2? It moves to the right eight units. It moves to the left four units. It moves to the left eight units. It moves to the right four units. im having a very hard time with this one

OpenStudy (anonymous):

sorry i have alot of questions

OpenStudy (triciaal):

the vertex is (h, k) y = a(x-h)^2 + k here a = 1 and k = 0 x-h = 0 x - - 4 sorry